5. (4pts each)Evaluate the following integrals. s sins (1) cos (1) dx a. dx b. sz * (ln(x))?
Answer #1

Solution [sin ()cos (1) ár Apply u-substitution u = sin (1) sin = 5ucos Substitute backu = sin (૬) - Sa () (5) s=9-0 dx= 3

Solution: : 1 x(In Apply u-substitution :u = = In(x) d (ln(x)) 1 x dx= xdu LE =fh Adjust integral boundaries: u=1n(x) Put in